# CoViS-Net: A Cooperative Visual Spatial Foundation Model for Multi-Robot Applications

### v1

@@ IV-B Model
We employ three models: $f_{\textrm{enc}}$, $f_{\textrm{pose}}$, and $f_{\textrm{agg}}$. The encoder $f_{\textrm{enc}}$ generates a common embedding of images, $f_{\textrm{pose}}$ performs a pairwise estimate of poses, and $f_{\textrm{agg}}$ spatially aggregates node features based on the output from $f_{\textrm{pose}}$. An overview of the model architecture is provided in Fig. 3.
@@
Transformers: Our methodology utilizes the Transformer architecture [57], distinguished by its self-attention mechanism. This mechanism computes the output as a weighted sum of the values ($\mathbf{V}$), with the weights determined by the compatibility of queries ($\mathbf{Q}$) and keys ($\mathbf{K}$) as per the formula: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}((\mathbf{QK}^{T})/\sqrt{d_{k}})\mathbf{V}$, where $d_{k}$ is the dimensionality of the keys. Typically, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are set to the same input features. Therefore, we refer to one attention block as $f_{\textrm{attn}}$ and to an instance of a multi-layer transformer, including both the attention mechanism and a position-wise feed-forward network, as $f_{\textrm{trans}}^{\textrm{inst}}$.
The Vision Transformer (ViT) [10] extends this architecture to images by dividing an image into patches treated as a sequence.
@@
Image encoder: We use DinoV2 [32] as a pre-trained encoder. DinoV2 is a vision foundation model based on ViT and pre-trained in a self-supervised manner on the LVD-142M dataset, which is a cumulation of various filtered datasets, including various versions of ImageNet [9] and Google Landmarks. We specifically chose DinoV2 due to its pre-training on a wide variety of images and demonstrated powerful feature representation as well as demonstrated depth understanding of scenes. We use the smallest distilled version of DinoV2 based on ViT-S/14 with $22.1$ M parameters for performance reasons. We refer to this model as $f_{\textrm{dino}}$.
For explicit communication, agents utilize a one-hop communication network, detailed further in Sec. V-C. Each node $v_{i}$ processes the image $I_{i}$ through the encoder function $f_{\textrm{enc}}$ to generate a node embedding $\mathbf{E}_{i}\in\mathbb{R}^{S\times F}$, which constitutes the sole information communicated to and utilized by neighboring nodes. Here, $S$ represents the sequence length (which can be reduced to adjust model performance, as explained in Sec. V), and $F$ denotes the size of each feature vector in the sequence.
The weights in $f_{\textrm{dino}}$ are frozen, and an instance of $f_{\textrm{trans}}^{\textrm{enc}}$, which is trained, is concatenated, resulting in $f_{\textrm{enc}}$ (Eq. 2).
@@
Freezing the weights of $f_{\textrm{dino}}$ facilitates the ability to use the raw output used independently of our proposed foundation model, meaning that a wide variety of open-source fine-tuned image detectors can be leveraged without running an additional forward pass.
Pairwise pose encoder: Our goal is to develop a pose prediction function, $f_{\textrm{pose}}$. At each root node $v_{i}$, we generate an image encoding $\mathbf{E}_{i}$ using $f_{\textrm{enc}}$. This encoding $\mathbf{E}_{i}$ is transmitted to neighboring nodes. Given the root and neighbor encodings $\mathbf{E}_{i}$ and $\mathbf{E}_{j}$, $f_{\textrm{pose}}$ estimates the pose of each neighbor relative to the root node. Specifically, for a given root node $v_{i}$ and each neighbor $v_{j}$, $f_{\textrm{pose}}$ calculates the set of estimated relative poses and uncertainties ${(\hat{\mathbf{p}}_{i,ij,,}:\hat{\mathbf{R}}_{i,ij},:\sigma^{2}_{ij}):\forall:v_{j}\in\mathcal{N}(v_{i})}$, with respect to $v_{i}$ in the coordinate frame $\mathcal{F}_{i}$, as well as edge embeddings ${\mathbf{E}_{ij}:\forall:v_{j}\in\mathcal{N}(v_{i})}$.  [^1] Although image overlap is not a strict assumption, our uncertainty measure $\sigma^{2}_{ij}$ enables us to identify image pairs that are unsuitable for joint estimation due to lack of feature correspondence, thus allowing the application to react accordingly, e.g., prioritize rotational alignment over positional alignment to regain certainty.
@@
$\sigma^{2}_{ij}$ $=(\sigma^{2}_{\mathrm{p},ij},\>\sigma^{2}_{\mathrm{R},ij})$ (6)
$f_{\textrm{pose}}(\mathbf{E}_{i},\mathbf{E}_{j})$ $=(\hat{\mathbf{p}}_{i,ij},\hat{\mathbf{R}}_{i,ij},\sigma^{2}_{ij}).$
We concatenate the node features along the sequence dimension, add the positional embedding, run the transformer and then extract the first element of the sequence dimension (Eq. 3), and eventually use this edge embedding to estimate poses and orientations (Eq. 4) and uncertainties (Eq. 5 and Eq. 6).
@@
IV-C Training
For training our models, we resort to supervised learning. This section describes the losses as well as our approach to estimating uncertainty.
Uncertainty estimation: In general, two kinds of uncertainties can be modeled. While epistemic uncertainty captures uncertainty in the model, i.e. uncertainty that can be improved by training the model on more data, aleatoric uncertainty captures noise inherent to a specific observation [17]. We argue that for estimating poses and applications to pose control in a robotic system, aleatoric uncertainty is much more important than epistemic uncertainty since a downstream application can act on the predicted aleatoric uncertainty to decrease it. Approaches such as Monte Carlo Dropout [13] or Model Ensembles [19] have previously used in similar applications [61], but these approaches model the epistemic uncertainty. Instead, we propose to use the Gaussian Negative Log Likelihood Loss, as introduced by Nix and Weigend [30] and revisited by Kendall and Gal [17]. As a generalization of the mean squared error (MSE) loss, it allows the model to not only predict a mean $\hat{\mu}$, but also a variance $\hat{\sigma}^{2}$, which is learned from data points $\mu$. The loss is defined as
$\mathcal{L}_{\text{GNLL}}(\mu,\hat{\mu},\hat{\sigma}^{2})=\frac{1}{2}\left(\log\left(\hat{\sigma}^{2}\right)+\frac{\left(\hat{\mu}-\mu\right)^{2}}{\hat{\sigma}^{2}}\right),$ (12)
where intuitively, the left component penalizes predictions of high variance, while the right component scales the MSE loss by the predicted variance to penalize errors more when the predicted uncertainty is low.
Estimating Rotations: Estimating rotation is inherently difficult due to discontinuities that occur in many orientation representations. In machine learning, quaternions are typically chosen to represent rotations due to their numerical efficiency and lack of singularities. Peretroukhin et al. [35] propose to represent rotations through a symmetric matrix that defines a Bingham distribution over unit quaternions, which allows them to predict epistemic uncertainty. To obtain an aleatoric uncertainty for the rotation, we combine Eq. 14 with Eq. 12 and let the model predict one variance for the rotation estimate, as shown in Eq. 15.
$d_{\text{quat}}\left(\hat{\mathbf{q}},\mathbf{q}\right)$ $=\min\left(\left\|\mathbf{q}-\hat{\mathbf{q}}\right\|_{2},\left\|\mathbf{q}+\hat{\mathbf{q}}\right\|_{2}\right)$ (13)
$\mathcal{L}_{\text{chord}}^{\text{GNLL}}\left(\mathbf{q},\hat{\mathbf{q}},\hat{\sigma}^{2}\right)$ $=\frac{1}{2}\left(\log\left(\hat{\sigma}^{2}\right)+\frac{\mathcal{L}_{\text{chord}}^{2}\left(\hat{\mathbf{q}},\mathbf{q}\right)}{\hat{\sigma}^{2}}\right)$ (15)
BEV prediction: For the BEV prediction, we use a combination of the Dice loss [52] $\mathcal{L}_{\mathrm{Dice}}$ and Binary Cross Entropy (BCE) Loss $\mathcal{L}_{\mathrm{BEC}}$ to achieve sharp and clear predictions. The BEV loss is defined as
$\mathcal{L}_{i,\mathrm{BEV}}=$ $\alpha\cdot\mathcal{L}_{\mathrm{Dice}}\left(\mathrm{BEV}_{i},\hat{\mathrm{BEV}_{i}}\right)$
$+(1-\alpha)\cdot\mathcal{L}_{\mathrm{BCE}}\left(\mathrm{BEV}_{i},\hat{\mathrm{BEV}_{i}}\right),$ (16)
where $0\leq\alpha\leq 1$ balances between the Dice and BCE objectives.
Training: We define the pose loss as
$\mathcal{L}_{i,ij,\mathrm{Pose}}=$ $\mathcal{L}_{\text{GNLL}}(\mathbf{p}_{i,ij},\hat{\mathbf{p}}_{i,ij},\sigma^{2}_{\mathrm{p},ij})$
$+\beta\>\mathcal{L}_{\text{chord}}^{\text{GNLL}}\left(\mathbf{R}_{i,ij},\hat{\mathbf{R}}_{i,ij},\sigma^{2}_{\mathrm{R},ij}\right),$ (17)
where $0\leq\beta\leq 1$ balances between the position and orientation loss, and, lastly, the total loss over the graph as
We find that the uncertainty terms in the loss are necessary for training, as many samples within the dataset do not contain overlaps, and are sometimes in two different (yet adjacent) rooms. Otherwise, the error from these difficult samples would dwarf the error from other samples, hampering learning.

### v2

@@ 4.1 Model
We employ four primary models: The encoder $f_{\textrm{enc}}$ generates a common embedding of images, $f_{\textrm{pose}}$ estimates pairwise poses, $f_{\textrm{agg}}$ aggregates node features, and $f_{\textrm{BEV}}$ predicts a two-dimensional BEV representation. An overview of the model architecture is provided in Fig. 3. Our model uses transformers [50, 51] in $f_{\textrm{enc}}$, $f_{\textrm{pose}}$ and $f_{\textrm{agg}}$. We provide more details in Sec. A.1.
Image Encoder: We use the smallest distilled version of DinoV2 [52] (ViT-S/14) as image encoder due to its robust feature representation and spatial understanding. Each robot $v_{i}$ processes the image $I_{i}$ through $f_{\textrm{enc}}$ to generate the communicated node embedding $\mathbf{E}_{i}\in\mathbb{R}^{S\times F}$, where $S$ represents the sequence length, and $F$ the feature vector size. We freeze the weights of DinoV2, enabling the use of various open-source downstream fine-tuned image detectors without an additional forward pass.
Pairwise pose encoder: For each ego-robot (i.e., root node $v_{i}$), the image encoding $\mathbf{E}_{i}$ is transmitted to neighboring robots. Given encodings $\mathbf{E}_{i}$ and $\mathbf{E}_{j}$, $f_{\textrm{pose}}$ estimates relative poses and uncertainties ${(\hat{\mathbf{p}}_{i,ij},\>\sigma^{2}_{\mathrm{p},ij},\>\hat{\mathbf{R}}_{i,ij},\>\sigma^{2}_{\mathrm{R},ij})}$ and edge embeddings ${\mathbf{E}_{ij}}$. For each incoming $\mathbf{E}_{j}\>\forall\>v_{j}\in\mathcal{N}(v_{i})$, we concatenate the node features along the sequence dimension, add the positional embedding, run the transformer and then extract the first element of the sequence dimension to obtain the edge embedding. We then use this edge embedding to estimate position, rotation and uncertainties, resulting in $f_{\textrm{pose}}(\mathbf{E}_{i},\mathbf{E}_{j})=(\hat{\mathbf{p}}_{i,ij},\>\sigma^{2}_{\mathrm{p},ij},\>\hat{\mathbf{R}}_{i,ij},\>\sigma^{2}_{\mathrm{R},ij})$.
Multi-Node Aggregation: We perform position-aware feature aggregation using node embeddings $\mathbf{E}_{j}$ from $f_{\textrm{pose}}$. The function $f_{\textrm{agg}}$ integrates information from all nodes $v_{j}\in\mathcal{N}(v_{i})\cup{v_{i}}$, producing a feature vector $\mathbf{F}_{i}$ for each node $v_{i}$, encapsulating the aggregated information from its local vicinity, which can be used for any downstream task in a transfer-learning context [43]. We concatenate node features along the sequence dimension and add the positional embedding. We then concatenate a nonlinear transformation of the edge features along the feature dimension. This operation is performed for all $v_{j}\in\mathcal{N}(v_{i})$, including self-loops to handle cases where $\mathcal{N}(v_{i})$ is empty by setting the edge embedding to the zero tensor. We aggregate over the neighbors and the ego-robot to produce the output embedding $\mathbf{F}_{i}$ from all edge embeddings $\mathbf{F}_{ij}$, which is used for downstream tasks, such as estimating a BEV representation $\hat{\mathrm{BEV}}_{i}=f_{\textrm{BEV}}(\mathbf{F}_{i})$.
4.2 Training
For training our models, we use supervised learning. This section describes the dataset, loss functions, and uncertainty estimation.
Dataset: We use the Habitat simulator [53, 54] and the HM3D dataset [55], consisting of $800$ scenes of 3D-scanned multi-floor buildings.The data is divided into training $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$ ($80\%$), testing $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ ($19\%$), and validation $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ ($1\%$) sets. We provide more details in Sec. A.2.
Uncertainty Estimation: Two types of uncertainties can be modeled: epistemic uncertainty, which captures uncertainty in the model and can be reduced with more data, and aleatoric uncertainty, which captures noise inherent to specific observations [56]. For pose estimation and pose control in robotic systems, aleatoric uncertainty is more important since downstream applications can act on it to reduce it. Traditional methods like Monte Carlo Dropout [57] and Model Ensembles [58] model epistemic uncertainty. Instead, we propose using the Gaussian Negative Log Likelihood (GNLL) Loss [59, 56], a generalization of the mean squared error (MSE) loss, allowing the model to predict both a mean $\hat{\mu}$ and a variance $\hat{\sigma}^{2}$, learned from data points $\mu$. The loss is defined as
$\mathcal{L}^{\text{GNLL}}(\mu,\hat{\mu},\hat{\sigma}^{2})=\frac{1}{2}\left(\log\left(\hat{\sigma}^{2}\right)+\frac{\left(\hat{\mu}-\mu\right)^{2}}{\hat{\sigma}^{2}}\right).$ (1)
The GNLL loss penalizes high variance predictions and scales the MSE loss by the predicted variance, penalizing errors more when predicted uncertainty is high.
Estimating Rotations: Estimating rotation is difficult due to discontinuities in many standard orientation representations. Quaternions are often used for their numerical efficiency and lack of singularities. Related work [60] represents rotations with a symmetric matrix defining a Bingham distribution over unit quaternions to predict epistemic uncertainty. We combine the chordal loss and GNLL to predict aleatoric uncertainty for rotations, as shown in the following equation:
$d_{\text{quat}}(\hat{\mathbf{q}},\mathbf{q})$ $=\min\left(\left\|\mathbf{q}-\hat{\mathbf{q}}\right\|_{2},\left\|\mathbf{q}+\hat{\mathbf{q}}\right\|_{2}\right)$ (2)
$\mathcal{L}_{\text{chord}}^{\text{GNLL}}\left(\mathbf{q},\hat{\mathbf{q}},\hat{\sigma}^{2}\right)$ $=\frac{1}{2}\left(\log\left(\hat{\sigma}^{2}\right)+\frac{\mathcal{L}_{\text{chord}}^{2}\left(\hat{\mathbf{q}},\mathbf{q}\right)}{\hat{\sigma}^{2}}\right).$ (4)
Training: We use a combination of the Dice loss [61] $\mathcal{L}_{\mathrm{Dice}}$ and Binary Cross Entropy (BCE) Loss $\mathcal{L}_{\mathrm{BEC}}$ weighted by the parameter $0\leq\alpha\leq 1$. This is commonly referred to as combo loss [62] and is typically utilized for slight class imbalance. We define the total loss as
$\mathcal{L}_{i,\mathrm{BEV}}$ $=\alpha\cdot\mathcal{L}_{\mathrm{Dice}}\left(\mathrm{BEV}_{i},\hat{\mathrm{BEV}}_{i}\right)+(1-\alpha)\cdot\mathcal{L}_{\mathrm{BCE}}\left(\mathrm{BEV}_{i},\hat{\mathrm{BEV}}_{i}\right)$ (5)
$\mathcal{L}_{i,ij,\mathrm{Pose}}$ $=\left(1-\beta\right)\mathcal{L}^{\text{GNLL}}(\mathbf{p}_{i,ij},\hat{\mathbf{p}}_{i,ij},\sigma^{2}_{\mathrm{p},ij})+\beta\>\mathcal{L}_{\text{chord}}^{\text{GNLL}}\left(\mathbf{R}_{i,ij},\hat{\mathbf{R}}_{i,ij},\sigma^{2}_{\mathrm{R},ij}\right)$ (6)
The parameter $0\leq\beta\leq 1$ balances between the position and orientation loss. Uncertainty terms are necessary for training, as many dataset samples lack overlap and may be in adjacent rooms, leading to significant errors from difficult samples that could hinder learning. We train on simulated data $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$ and provide further details in Sec. A.4.
5 Experiments and Results
We demonstrate the performance of our trained models on the test set $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ and custom real-world test sets $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$. First, we outline the metrics used. Then, we introduce the real-world dataset. Next, we analyze model performance on both simulation and real-world data. Finally, we compile and deploy our model in a real-world, real-time trajectory tracking task.
Metrics: We compute the Euclidean distance between the predicted and ground truth poses as $D_{\mathrm{pos}}(\mathbf{p}_{i,ij},\hat{\mathbf{p}}_{i,ij})$ and the geodesic distance between rotations as $D_{\mathrm{rot}}(\mathbf{R}_{i,ij},\hat{\mathbf{R}}_{i,ij})$ for all edges and report median errors. We categorize the median pose error into four categories. All includes all edges. Visible and Invisible edges are determined by FOV overlap (an edge is invisible if $D_{\mathrm{rot}}(\mathbf{R}_{i,ij},\mathbf{0})>\mathrm{FOV}$). Invisible Filtered samples use the predicted position uncertainty $f_{\textrm{p}}^{\sigma}{ij}$ to reject high-uncertainty pose estimations based on the Youden’s index [63]. For simulation results, we also report the Dice and IoU metric. We provide more details in Sec. B.1.
5.1 Dataset Evaluation
We conducted experiments across various model variants trained with different parameters, reporting on our model’s efficiency. We present quantitative results in Tab. 1 and qualitative findings in Fig. 2.
Real-World Dataset: Our setup includes four Cambridge RoboMasters [64] and one Unitree Go1 Quadruped, each equipped with local compute, a monocular camera and WiFi dongles for adhoc networking. Our custom real-world test dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ covers a diverse and challenging range of indoor and outdoor scenes. We provide more details in Sec. A.3.
Table 1: Ablation study over the number of patches $S$ and size of features $F$ per patch. We report the BEV representation performance and the median error for poses on the dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ and $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
Model $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$
S F Dice IoU All Invis. Filt. Invisible Visible
256 48 69.1 57.1 36 cm 8.3° 61 cm 6.8° 97 cm 7.9° 33 cm 5.8°
128 96 68.8 56.8 40 cm 8.4° 55 cm 7.7° 97 cm 7.4° 32 cm 5.6°
128 48 67.9 56.1 38 cm 7.7° 67 cm 9.9° 113 cm 9.6° 29 cm 5.7°
128 24 66.7 54.6 50 cm 9.5° 83 cm 11.2° 112 cm 9.7° 31 cm 5.7°
64 48 61.0 43.5 51 cm 9.6° 81 cm 9.4° 119 cm 10.8° 36 cm 6.3°
1 3072 47.0 1.4 144 cm 89.9° 123 cm 25.7° 122 cm 25.8° 93 cm 138.2°
1 348 47.1 1.4 84 cm 11.7° 150 cm 164.1° 135 cm 37.9° 80 cm 11.1°
Results: Tab. 1 details the performance across seven different experiments. We evaluated the impact of image embedding sizes $\mathbf{E}_{i}$ on communication efficiency. The IoU and Dice metrics for the BEV representation, alongside median pose estimates for all samples in the dataset, are provided. Experiments show that the sequence length $S$ significantly influences BEV representation and pose estimate performance more than feature size $F$. We set $S$ to 1 for two experiments: (1) increasing $F$ to match the size of the embeddings for other experiments, and (2) decreasing $F$, effectively reducing the transformer to a CNN-only architecture by removing the attention mechanism. Both experiments result in reduced performance, highlighting the importance of attention in our model. The best-performing model with $S=256$ achieves an IoU of $0.571$, a Dice score of $0.691$, and median pose accuracy of $36$ cm and $8.3^{\circ}$, which is roughly equal to one robot body length. In contrast, models with $S=1$ show significant degradation, with nearly unusable BEV representation predictions and pose errors on the scale of meters. Models with $F=24$ and $S=64$ performed slightly worse than those with larger $F$ or $S$. We provide additional results in Sec. B.2. Evaluation on the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ focuses on position and rotation accuracy, excluding quantitative BEV performance due to unavailable ground truth data. The results align with $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$, showing a correlation between the number of patches $S$ and features $F$ and pose estimation performance. The rotation estimates do not significantly differ between visible and invisible samples, demonstrating robust pose estimation without requiring visual overlap. The most accurate model achieves a median localization error of 97 cm and $7.9^{\circ}$ for invisible samples, and 33 cm and $5.8^{\circ}$ for visible samples. Note that $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ samples are at most 2 m apart, resulting in a worst-case error of up to 4 m. These findings demonstrate our model’s capability to reliably estimate poses, even with minimal overlap, and accurately predict higher variances under adverse conditions. We attribute this performance to our novel transformer-based pose prediction architecture, as well as the robust spational representation capabilities of the DinoV2 encoder. The model goes beyond mere feature matching, achieving sophisticated scene understanding that allows it to “imagine” unseen portions of the scene based on learned data priors, thus facilitating pose estimation. We provide additional results in Sec. B.3.
Finally, we conduct a baseline comparison with two feature-matching-based approaches, as summarized in Table 2. Both baseline methods extract and match features across two images to estimate the essential matrix, deriving rotation and scale-less translation vectors. The first baseline uses OpenCV for ORB feature extraction followed by brute-force matching. The second baseline, LightGlue, uses a neural network-based approach for feature extraction and matching. Given both baselines estimate scale-less transformations, we adopt the same metric used in LightGlue [39], computing the AUC at thresholds of 20, 45, and 90 degrees based on the maximum error in rotation and translation. Our findings indicate that our model surpasses both baselines in performance across image pairs with and without visual overlap. Notably, while both baselines require visual overlap to function effectively, our approach does not, highlighting its robustness and applicability in a broader range of scenarios.
Table 2: We report the AUC metric at 20, 45 and 90°on a first-principles and an NN-based baseline.
All Invisible Visible
AUC@ 20 45 90 20 45 90 20 45 90
ORB/OpenCV [65] 6.45 15.43 28.26 0.16 1.60 7.43 9.41 21.92 38.03
LightGlue [39] 16.30 27.14 37.63 0.02 0.09 0.31 23.94 39.82 55.14
Ours 25.23 47.63 66.34 9.22 24.40 45.74 32.74 58.53 76.01
5.2 Real-World Pose Control
In this section, we describe our experiment setup (including model deployment, wireless communication, and controllers), demonstrating real-world multi-robot control tasks.
Figure 4: We evaluate the tracking performance of our model and uncertainty-aware controller on reference trajectories with two follower robots (blue and orange) positioned to the left and right of the leader robot (green). We report tracking performance over time for position and rotation, as well as the predicted uncertainties. The reference trajectory is a Figure Eight, with the leader always facing the direction of movement. We show the trajectory for 120 s and the tracking error for the first 60 s. The tracking performance is consistent across multiple runs.
Model deployment: We deploy a model with $F=24$ and $S=128$. After training in PyTorch, we compile $f_{\textrm{enc}}$, $f_{\textrm{pose}}$, and $f_{\textrm{agg}}$ for decentralized deployment. Using TensorRT, we achieve sub-30 ms processing times. We provide a more details and a runtime analysis in Sec. B.4.
Ad-hoc WiFi: Image embeddings are sent between robots at 15 Hz, each over 6 KiB. The system’s distributed nature suits broadcast communication, allowing scalable operation. However, IEEE 802.11/WiFi’s fallback to low bit rates for broadcast increases frame loss probability, worsening with more robots. Higher-level systems like ROS2’s CycloneDDS manage retransmissions but can’t detect network overloads, leading to increased packet loss. We design a custom protocol that dynamically adjusts messaging load based on network conditions and uses TDMA scheduling to maximize frame reception without compromising data rates. We provide more details in Sec. B.4.
Experiment: We conduct qualitative experiments in various environments, including outdoors, and quantitative experiments. Two follower robots maintain a fixed orientation and distance from a leader robot on a predefined trajectory. A PD controller manages relative pose control, adjusting based on predicted uncertainty and selectively deactivating if uncertainty thresholds are exceeded. This ensures that even with unreliable estimates, robots maintain orientation toward the leader until reliable estimates are available. We assess tracking accuracy by comparing ground truth and estimated poses, focusing on position and rotation errors and predicted uncertainties.
Results: Fig. 5 shows snapshots from four different real-world deployments, and Fig. 4 shows the trajectories of all robots for all quantitative experiments, as well as a visualization of the position and rotation error over time. We provide additional results in Sec. B.4.
0s1s2s3s4s5s
[…]
The qualitative experiments show a robust system that generalizes to a wide range of scenarios, including outdoors, albeit with a scale prediction that is less reliable outdoors. Our model was trained exclusively on indoor data and generalizes to outdoor data due to the wide range of training data used on the DinoV2 encoder. We furthermore show platform agnosticism by deploying our model to a heterogeneous team of robots, lead by a quadruped (second row in Fig. 5).
In quantitative experiments, the tracking error oscillates throughout the trajectories as the followers reactively adjust their positions to the leader’s movements. Follower A starts facing the opposite direction of the leader, resulting in high position variance and low rotation variance, which decreases as the robot aligns with the leader. An erroneous rotation prediction is accompanied by high uncertainty. As the leader moves, one follower is inside and the other outside the trajectory, switching roles depending on the section of the figure. We report a median tracking error of 48 cm and 5.3° for Follower A, and 26 cm and 5.4° for Follower B (corresponding roughly to one robot body length).
6 Limitations and Conclusion
In this work, we introduced a first-of-its-kind cooperative visual spatial foundation model that leverages existing large vision models, and demonstrated its use in multi-robot control. From monocular camera images only, our model is able to output accurate relative pose estimates and BEV representations by aggregating information across neighboring robots. We validated our model on a custom real-world test set and demonstrated real-time multi-robot control. Our model accurately predicts poses, even without pixel-level correspondences, and correctly predicts BEV representations for obscured regions. In further real-world tests, we demonstrated accurate trajectory tracking, paving the way for more complex vision-only robot tasks. Our model estimates relative poses with a median error of 33 cm and 5.8° for visible edges, and 97 cm and 5.9° for invisible edges, roughly equivalent to one body length of our robot. In future work, we plan to include temporal data, apply this model to other downstream tasks and combine it with learning-based multi-agent control policies.
Acknowledgments
This work was supported in part by European Research Council (ERC) Project 949949 (gAIa) and in part by ARL DCIST CRA W911NF-17-2-0181. J. Blumenkamp acknowledges the support of the ‘Studienstiftung des deutschen Volkes’ and an EPSRC tuition fee grant. We also acknowledge a gift from Arm.
Appendix A Model Training
A.1 Model
@@
Our methodology utilizes the Transformer architecture [50]. Self-attention computes the output as a weighted sum of the values ($\mathbf{V}$), with the weights determined by the compatibility of queries ($\mathbf{Q}$) and keys ($\mathbf{K}$) as per the formula: $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}((\mathbf{QK}^{T})/\sqrt{d_{k}})\mathbf{V}$, where $d_{k}$ is the dimensionality of the keys. Typically, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are set to the same input features.
@@
DinoV2 is a vision foundation model based on the vision transformer (ViT) and pre-trained in a self-supervised manner on the LVD-142M dataset, which is a cumulation of various filtered datasets, including various versions of ImageNet [66] and Google Landmarks.
The encoder $f_{\textrm{enc}}$ consists of the pre-trained DinoV2 vision transformer and two more transformer layers previously described as $f_{\textrm{trans}}^{\textrm{enc}}$. We use the smallest distilled version of DinoV2 based on ViT-S/14 with $22.1$ M parameters for performance reasons. The input feature size is $224\times 224$, of which DinoV2 generates an embedding of size $S=256$ and $F=384$. We add a linear layer that transforms $F$ down to the parameter described in this paper (e.g. $24$) while maintaining $S$. The output of this is then cropped to the $S$ described in this paper (e.g. 128) by cropping the sequence length at that position.
@@
The pairwise pose encoder $f_{\textrm{pose}}$ consists of transformer $f_{\textrm{trans}}^{\textrm{pose}}$, which is assembled from five transformer blocks, and four MLPs $f^{\mu}_{\textrm{p}}$, $f_{\textrm{p}}^{\sigma}$, $f^{\mu}_{\textrm{R}}$, $f_{\textrm{R}}^{\sigma}$.
@@
$f_{\textrm{pose}}(\mathbf{E}_{i},\mathbf{E}_{j})$ $=(\hat{\mathbf{p}}_{i,ij},\hat{\mathbf{R}}_{i,ij},\sigma^{2}_{\mathrm{p},ij},\sigma^{2}_{\mathrm{R},ij}).$
We concatenate the node features along the sequence dimension, add the positional embedding, run the transformer and then extract the first element of the sequence dimension (Eq. 10), and eventually use this edge embedding to estimate poses and orientations (Eq. 11) and uncertainties (Eq. 12). We first project the embedding $\mathbf{E}_{i}$ to a larger feature size $F=192$ using a linear layer. Each transformer layer has $F=192$ in and out features. The learnable transformer positional embedding $\mathbf{S}_{\mathrm{pose}}$ is initialized to match $F$ and $S$, with normal distribution initialization of $\sigma=0.02$. The MLPs $f^{\mu}_{\textrm{p}}$, $f_{\textrm{p}}^{\sigma}$, $f^{\mu}_{\textrm{R}}$, $f_{\textrm{R}}^{\sigma}$ are represented by one linear layer mapping the output of the transformer from $F=192$ to $F=17$ (position: 3, position uncertainty: 3, orientation: 10, orientation uncertainty: 1).
@@
A.2 Training Dataset
We utilize the Habitat simulator [53, 54] along with the training set of the HM3D dataset [55], which comprises $800$ scenes of 3D-scanned real-world multi-floor buildings. We select Habitat and HM3D for their photorealism and the availability of ground-truth pose information, as well as detailed navigation mesh information that can be converted to a BEV representation of the scene. We instantiate a single agent equipped with a calibrated camera with a field of view of $120^{\circ}$. For each scene, we randomly sample several uniform agent positions based on the total area of the scene. For each sample, the camera orientation is randomized across all three axes. Additionally, for each scene, we extract the processed navigation mesh, which indicates areas that are navigable by a mobile robot, or not navigable due to obstacles or overhangs. We extract $3,816,288$ images from a calibrated camera with a $120^{\circ}$ field of view from $800$ scenes and $1,411$ floors, along with corresponding BEV representations.
The samples of the datasets $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$ $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ and $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ are randomly sampled to be within orientation range $[-\pi,\pi]$ for roll and pitch, and within a height of 0.0 to 2.0 meters, or less if the ceiling is at a lower height, resulting in uniformly distributed 6D poses.
We divide our data into training $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$ ($80\%$), testing $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ ($19\%$), and validation $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ ($1\%$) sets. From these sets, we sample pairs of $N_{\mathrm{max}}=5$ images and poses by (1) uniformly sampling $25\%$ of individual images from each dataset, and then (2) uniformly sampling $N_{\mathrm{max}}-1$ images within a distance of $d_{\mathrm{max}}=2\textrm{ m}$ around each sample from (1). This process yields $763,257$ training and $276,680$ test tuples, each of which are at most $2d_{\mathrm{max}}$ apart and pointing in random directions, thus providing realistically close and distant data samples for estimating both poses and uncertainties in case of no overlaps.
A.3 Validation Dataset
Data collection involves a base station with dual cameras and a 2D Lidar, plus a distinctively shaped marker on each robot. We manually navigated robots around the base station to capture a variety of relative poses, then used a custom tool to label Lidar data for pose extraction.
All robots are equipped with an NVidia Jetson Orin NX 16GB and a forward-facing camera, rectified to a $120^{\circ}$ FOV with an image resolution of $224\times 224$ pixels. We collected the real-world test set $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ in eight indoor scenes. It consists of 5692 images. We sample image pairs to ensure a uniform distribution of relative positions from 0 m to 2 m. This results in 14008 sets of samples, each with three images, totaling 84048 pose edges, $32\%$ which have no visual overlap. Our dataset covers a diverse and challenging range of indoor scenes, as well as one outdoor scene.
We explain the setup of our data collection unit in Fig. 6. We visualize the five different scenes we collected the dataset from as well as the data distribution from each respective scene in Fig. 7.
Figure 7: We collect the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ from eight unique scenes, showing different challenges, from cluttered environments over scenes with high ceilings to sunny floors. The first row shows a picture summary of the scene, taken from first-person-view perspective. The second to fourth row shows three samples from the real-world dataset, from the perspective of the robot. The fifth row shows the distribution of positions for each scene and the sixth row shows the distribution of relative angles between nodes.
A.4 Training